OpenStudy (jiteshmeghwal9):

\The total no. of integral solutions of equation \(x+y+z+t=29\) where \[x \ge 1, y \ge 2, z \ge 3,t \ge 0\]

Parth (parthkohli):

Let \(x = a+1\), \(y = b+2\), \(z = c+3\) where \(a \ge 0, b\ge 0, c\ge0\). Now we have\[a+1+b+2+c+3+t=29\]\[\Rightarrow a + b + c + t = 23\]where all variables are nonnegative

Parth (parthkohli):

\[\binom{23+4-1}{4-1}\]Accordingly, this is the answer.

OpenStudy (jiteshmeghwal9):

yes 

OpenStudy (spring98):

wowza that is hard

OpenStudy (empty):

That was a clever trick to make all the variables the same like that, cool.

OpenStudy (zzr0ck3r):

I will name my first born Parth